### v1

@@ 1.2 Main results and conjectures
The COMB strategy is asymptotically optimal only for $n\leq 4$ experts.
@@
The non-COMB strategy $(1,0,1,0,0)$ is the only asymptotically optimal strategy for $n=5$ experts.
@@
There are no asymptotically optimal strategies for $n\geq 6$ experts.
@@
It is important to point out that there are certainly some limitations to our numerical results. In particular, we cannot solve the PDE (1.1) numerically on the full space $\mathbb{R}^{n}$, and must restrict our attention to a compact subset. Our numerical results are obtained over the box $[-1,1]^{n}$. We cannot rule out, for example, that Conjecture 1.2 fails somewhere outside of the box $[-1,1]^{n}$. However, we do find that for $n=2,3,4$ experts, the optimality observed on the box matches perfectly with the global theory. The negative results of Conjectures 1.1 and 1.3 do not suffer the same limitations, since non-optimality need only be observed at a single point.
@@ 2 Analysis of a general numerical scheme
The class of monotone equations of the form (2.6) is closely related to the wide stencil finite difference schemes introduced and studied by Oberman [37]. The main difference in this section is that we are focused on the unbounded domain $\mathbb{Z}^{n}_{h}$, and we are interested in properties of the solution $u$, such as Lipschitzness, convexity, permutation invariance, etc., that hold under certain structure conditions and the source term $g$.
@@ 2.1 Existence and uniqueness
$\leq\frac{1}{h^{2}}(v(x_{0}+h\mathbf{v})-u(x_{0})+v(x_{0}-h\mathbf{v})-v(x_{0}))=\nabla^{2}_{h}v(x_{0},\mathbf{v})$
@@
since $h\leq 1$ and $|\mathbf{v}|\geq 1$. Since $F(0)=0$ we have
@@ 2.3 Convergence rates
Existence and uniqueness of a linear growth viscosity solution to (2.21) is standard material for viscosity solutions, and proofs can be found in [18, 9]. For use later on, we recall the comparison principle for (2.21) in Lemma 2.12 below. The proof of this result is standard in viscosity solution theory; a self-contained proof can be found in [9, Lemma 12.17].
@@
Assume $F$ is monotone, has width $N$, and satisfies $F(0)=0$. Let $u\in\text{USC}(\mathbb{R}^{n})$ be a viscosity subsolution of (2.21) and let $v\in\text{LSC}(\mathbb{R}^{n})$ be a viscosity supersolution of (2.21). If
@@
Using this comparison principle and the Perron method, we can prove existence of a linear growth solution (Theorem 2.13 below). The application of the Perron method is standard and a self-contained proof can be found in [9, Theorem 12.18]. [^2] In fact, the proofs do not require Lipschitzness; uniform continuity is sufficient.
@@ 3.2 Reducing the domain to a sector
As a final remark, the computational costs are more expensive than the memory required to store the grid, since we must compute the maximum over $\mathbf{v}\in{\mathcal{B}}^{n}$ in the operator, which is a maximum over $2^{n}$ directions. Thus, the computational cost admits an additional $2^{n}$ factor over the memory costs.
@@ 5 Conclusion and future work
We expect these numerical methods could be extended to a few more experts, perhaps the $n=11$ and $n=12$ expect problems, using parallel processing or computational clusters with vastly more memory. The finite horizon problem should be amenable to similar techniques. There are also other prediction with expert advice PDEs, in particular the history-dependent experts setting [11, 20], that would benefit from numerical explorations. In terms of theory, we posed a number of conjectures and open problems that stem from this work in Section 1.2 that would be interesting to explore in future work.

### v2

@@ 1.2 Main results and conjectures
The COMB strategy is globally asymptotically optimal (i.e., on the sector $\mathbb{S}_{n}$) only for $n\leq 4$ experts.
@@
The non-COMB strategy $(1,0,1,0,0)$ is the only globally asymptotically optimal strategy for $n=5$ experts.
@@
There is no globally asymptotically optimal adversary strategy that is constant on the sector $\mathbb{S}_{n}$ for $n\geq 6$ experts.
@@
It is important to point out that there are certainly some limitations to our numerical results. In particular, we cannot solve the PDE (1.1) numerically on the full space $\mathbb{R}^{n}$, and must restrict our attention to a compact subset. Our numerical results are obtained over the box $[-1,1]^{n}$. We cannot rule out, for example, that Conjecture 1.2 fails somewhere outside of the box $[-1,1]^{n}$. However, we do find that for $n=2,3,4$ experts, the optimality observed on the box matches perfectly with the global theory. The negative results of Conjectures 1.1 and 1.3 do not suffer the same limitations, since non-optimality need only be observed at a single point. Finally, our numerical convergence rates rely on classical regularity of the viscosity solution, which is only known for $n\leq 4$ experts.
Remark 1.5.
We also remark that it may be possible to use our techniques for reducing the computational grid to the sector $\mathbb{S}_{n}$ in combination with the dynamic programming approaches of [27, 17]. We leave this interesting direction to future work.
@@ 2 Analysis of a general numerical scheme
We note that $X,Y:{\mathcal{H}}_{n}\to\mathbb{R}$ are real-valued functions, so $X\leq Y$ means that $X(\mathbf{v})\leq Y(\mathbf{v})$ for all $\mathbf{v}\in\mathbb{Z}_{n}$. The class of monotone equations of the form (2.6) is closely related to the wide stencil finite difference schemes introduced and studied by Oberman [38]. The main difference in this section is that we are focused on the unbounded domain $\mathbb{Z}^{n}_{h}$, and we are interested in properties of the solution $u$, such as Lipschitzness, convexity, permutation invariance, etc., that hold under certain structure conditions and the source term $g$.
@@ 2.1 Existence and uniqueness
$\leq\frac{1}{h^{2}}(v(x_{0}+h\mathbf{v})-v(x_{0})+v(x_{0}-h\mathbf{v})-v(x_{0}))=\nabla^{2}_{h}v(x_{0},\mathbf{v})$
@@
where we used a Taylor series expansion of $\nabla^{2}_{h}\psi(x,\mathbf{v})$ in the last line, along with $h\leq 1$ and $|\mathbf{v}|\geq 1$. Since $F(0)=0$ we have
@@ 2.3 Convergence rates
Existence and uniqueness of a linear growth viscosity solution to (2.21) is standard material for viscosity solutions, and proofs can be found in [19, 10]. For use later on, we recall the comparison principle for (2.21) in Lemma 2.12 below. The proof of this result is standard in viscosity solution theory; a self-contained proof can be found in [10, Lemma 12.17].
@@
Assume $F$ is uniformly continuous, monotone, has width $N$, and satisfies $F(0)=0$. Let $u\in\text{USC}(\mathbb{R}^{n})$ be a viscosity subsolution of (2.21) and let $v\in\text{LSC}(\mathbb{R}^{n})$ be a viscosity supersolution of (2.21). If
@@
Using this comparison principle and the Perron method, we can prove existence of a linear growth solution (Theorem 2.13 below). The application of the Perron method is standard and a self-contained proof can be found in [10, Theorem 12.18]. [^2] In fact, the proofs do not require Lipschitzness; uniform continuity is sufficient.
@@ 3.2 Reducing the domain to a sector
Remark 3.5.
It is important to point out that the computational costs are more expensive than the memory required to store the grid, since we must compute the maximum over $\mathbf{v}\in{\mathcal{B}}^{n}$ in the operator, which is a maximum over $2^{n}$ directions. Thus, the computational cost admits an additional $2^{n}$ factor over the memory costs.
@@ 5 Conclusion and future work
We expect these numerical methods could be extended to a few more experts, perhaps the $n=11$ and $n=12$ expert problems, using parallel processing or computational clusters with vastly more memory. The finite horizon problem should be amenable to similar techniques. There are also other prediction with expert advice PDEs, in particular the history-dependent experts setting [12, 21], that would benefit from numerical explorations. In terms of theory, we posed a number of conjectures and open problems that stem from this work in Section 1.2 that would be interesting to explore in future work.